# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 12. Vertex-transitive graphs
We may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$. By 9.41$(U,W)$ is a totally-nested tetra-separation. By Lemma 12.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
We claim that $G[U]=K_{4}$. Assume otherwise, so $G[W]=K_{4}$. Then $G[W]=G_{s}$ for some bag $G_{s}\neq G_{t}$. Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma 9.14. But then $\alpha(\mathcal{O})<4$ contradicts Lemma 9.51.
Hence $G[U]=K_{4}$. Next, we claim that $S(U,W)$ contains no inward edge, and assume for a contradiction that $x_{1}\in S(U,W)$ is an inward edge with endvertex $u$ in the interior of $G_{t}$. By Lemma 9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$. In particular, $y_{1}$ cannot be an outward edge. Since $S(U,W)$ consists of four edges and $y_{1}$ is one of them, $y_{1}$ must be an inward edge. Let $u$ be the endvertex of $y_{1}$ in the interior of $G_{t}$. Then $u\neq v$ by the matching-condition for $(U,W)$. Let $w$ be a neighbour of $x^{\prime}_{1}$ in $G$ but outside of $G_{t}$. Then the vertex-set $\{u,v,w\}$ is independent in $G$ and consists entirely of neighbours of $x^{\prime}_{1}$. Since $G[U]=K_{4}$ and $S(U,W)$ consists of four edges, $G$ is 4-regular and every vertex lies in a $K_{4}$ by vertex-transitivity. But $\{u,v,w\}$ witnesses that $x^{\prime}_{1}$ does not lie in a $K_{4}$, a contradiction. Therefore, $S(U,W)$ contains no inward edge.
@@
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$. Write $O=t(0)\,t(1)\ldots t(m)\,t(0)$ with indices in $\mathbb{Z}_{m+1}$. By Lemma 12.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles. Then, by Lemma 9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected. Hence, both of these torsos are good as they neighbour torsos that are 4-cycles. By induction, we get that $O$ has even length, that bags $G_{t(i)}$ are $K_{4}$’s for even $i$, and that torsos $H_{t(i)}$ are 4-cycles for odd $i$. By Lemma 9.51, there are $\geqslant 4$ bags that are $K_{4}$’s. It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$. ∎
@@
Write $\mathcal{O}=:(O,\mathcal{G})$. By Lemma 12.12, all bags of $\mathcal{O}$ are bad. Let $H_{t}$ be a torso of $\mathcal{O}$ on five vertices. Since the unique interior vertex $x$ of $H_{t}$ has degree 4, the graph $G$ is 4-regular. Since $H_{t}$ is bad, it is neighboured by a 3-connected torso $H_{s}$. Since $H_{s}$ is bad as well, it has $\leqslant 5$ vertices. That is, the torso of $H_{s}$ either is a $K_{4}$ or has exactly five vertices.
@@
It follows by induction that all torsos of $\mathcal{O}$ have exactly five vertices. By Lemma 12.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag. By 4-regularity and since every bag looks like this, these are all the edges that there are in the bags.
Since each bag has a unique interior vertex, the set $X$ of all these vertices has size $|O|$. Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma 9.51. If $|O|\geqslant 5$, then every vertex in $X$ is contained in exactly two induced 4-cycles in $G$, while every vertex in an adhesion set is contained in exactly one induced 4-cycle in $G$, contradicting vertex-transitivity. Therefore, we have $|O|=4$. Consider the graph $H:=G-X$, which is either an 8-cycle or a disjoint union of two 4-cycles. If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in $V(G-X)$ is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity. Therefore $H$ is a disjoint union of two 4-cycles, and it follows that $G$ is the line graph of the cube. ∎
@@
If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma 12.12, and it is of length $\geqslant 8$ by Lemma 9.51. Hence, we may assume that all torsos of $\mathcal{O}$ are bad.
We claim that $\mathcal{O}$ has no 4-cycle as torso. Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma 9.4, and they are good, a contradiction. Thus,
@@
We claim that if some torso of $\mathcal{O}$ is a triangle, then also its neighbouring torsos are triangles. For this, let $H_{t}$ and $H_{s}$ be neighbouring torsos with $H_{t}$ a triangle. By 8, the torso $H_{s}$ is either a triangle or a $K_{4}$. Assume for a contradiction that $H_{s}=K_{4}$. Let $\{u,v\}$ be the adhesion-set in between $H_{t}$ and $H_{s}$, so that $v$ is the tip of the triangle-torso $H_{t}$, say. Since every vertex has degree $\geqslant 4$, the edge $uv$ must exist in $G$, and $G$ must be 4-regular. Let $w$ be the vertex of the triangle $H_{t}$ besides $u$ and $v$. Let $G_{r}$ be the bag neighbouring $G_{t}$ besides $G_{s}$. Then $v$ has exactly one neighbour in the bag $G_{r}$. By 8, the torso $H_{r}$ can only be a triangle or a $K_{4}$. As $v$ has exactly one neighbour in the bag $G_{r}$, the torso $H_{r}$ cannot be a $K_{4}$, so it must be a triangle. But then the edge $vw$ exists by Lemma 9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity. Therefore, every triangle-torso of $\mathcal{O}$ is neighboured by two triangle-torsos.
So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma 9.4. Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma 9.51 and since $G\neq K_{5}$.
So we may assume that $\mathcal{O}$ has no triangle-torsos. Then by 8, all torsos of $\mathcal{O}$ are $K_{4}$’s. Then either all adhesion-sets are spanned by edges in $G$, in which case $G$ is a cycle of $K_{4}$-bags; or none of them are, in which case $G$ is a cycle of $K_{2,2}$-bags. In both cases, the length is $\geqslant 4$ by Lemma 9.51. ∎
@@
By the (Crossing Lemma). (4.1), all links have the same size 00, $1$ or $2$. If all links have size 00, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma 12.7. If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma 12.10. Otherwise all links have size $2$, and we apply Lemma 12.16. ∎
Observation 12.18.
Let $G^{\prime}$ be the expansion of a graph $G$. Let $v$ be a vertex of degree $d$ in $G$, and let $C_{v}$ be the clique of $G^{\prime}$ that corresponds to $v$. Then every vertex in $C_{v}$ has degree $d$ in $G^{\prime}$.∎
Observation 12.19.
Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the expansion of $G$. Then there is a unique partition $\mathcal{P}$ of $V(G^{\prime})$ into (the vertex-sets of) $|\mathcal{P}|=|V(G)|$ cliques. Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to $G$.∎
Observation 12.20.
The expansion of a graph $G$ is isomorphic to the graph $H$ with
$V(H)$ =${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
$E(H)$ =${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
@@
(i) $\Rightarrow$ (ii). Let $(u_{1},v_{1}),(u_{2},v_{2})$ be ordered pairs of vertices in $G$ with $u_{i}v_{i}\in E(G)$. We will find an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. For this, we abbreviate $x_{i}:=(u_{i},v_{i})$. Let $P_{i}\in\mathcal{P}$ be the unique partition-class of $\mathcal{P}$ that contains $x_{i}$. Since $G^{\prime}$ is vertex-transitive by assumption, there is an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. Note that $\varphi$ then maps $P_{1}$ to $P_{2}$. We define $\varphi:=\gamma\circ\varphi^{\prime}\circ\gamma^{-1}$ by abuse of notation, using that $\operatorname{Aut}(G^{\prime})$ acts on $\mathcal{P}$. Clearly, $\varphi$ is a bijection $V(G)\to V(G)$. Moreover, we have
[…]
@@
(ii) $\Rightarrow$ (i). We assume without loss of generality that $G^{\prime}$ takes the form described in 12.20. Let $x_{1},x_{2}\in V(G^{\prime})$ be two vertices. We will find an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. We write $x_{i}=:(u_{i},v_{i})$ using 12.20, so $u_{i}v_{i}$ is an edge of $G$. Since $G$ is arc-transitive, there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. Now $\varphi$ defines an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that sends every vertex $x=(u,v)$ to the vertex $\varphi^{\prime}(x):=(\varphi(u),\varphi(v))$. The automorphism $\varphi^{\prime}$ satisfies
@@
By Lemma 12.21, it suffices to show that $G$ is not arc-transitive. Indeed, if $u$ is a vertex in $A$, then $G[N(u)]=K_{1}\sqcup K_{r-1}$ with $r-1\geqslant 2$. Let $v_{1}$ be the isolated vertex in $G[N(u)]$ and let $v_{2}\neq v_{1}$ be another vertex in $N(u)$. Then there is no automorphism of $G$ that maps $(u,v_{1})$ to $(u,v_{2})$. ∎
@@
Follows from Corollary 12.22. ∎
@@
Let $G$ be a cycle of triangle-bags of length $\geqslant 6$, and let $\mathcal{O}$ be the corresponding cycle-decomposition. Note that every free edge of $\mathcal{O}$ is contained in exactly one triangle of $G$, while every other edge of $G$ is contained in exactly two triangles. So, $G$ is not edge-transitive. In particular, it is not arc-transitive. By Lemma 12.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive. ∎
@@
This follows from Corollary 12.17 together with the following observations. First, a cycle of $K_{4}$-bags of length $\geqslant 4$ is not 4-regular. Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary 12.22. Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary 12.24. ∎
@@
Let $G$ be a vertex-transitive essentially 5-connected graph. Then either $G$ is quasi-5-connected or there is a partition $\mathcal{P}$ of $V(G)$ into vertex-sets of $K_{4}$’s such that the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
@@
If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma 2.3 (i). Hence, we may assume that $G$ has a tetra-separation $(A,B)$. Since $G$ is essentially 5-connected, the separator of $(A,B)$ consists of four edges and there is a side, say $A$, that induces a $K_{4}$ in $G$. Let $u\in A$ and observe that $G[N(u)]=K_{1}\sqcup K_{3}$; in particular, $u$ is contained in exactly one $K_{4}$. By vertex-transitivity, so is every vertex of $G$. It follows that $G$ admits a partition of its vertex-set into the vertex-sets of $K_{4}$’s. Moreover, $G$ is 4-regular since $u$ has degree 4 in $G$. Then it follows that the edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. ∎
@@
Let $G$ be a vertex-transitive essentially 5-connected graph. Then $G$ is one of the following:
@@
Assume that $G$ is not quasi-5-connected. Then by Lemma 12.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. We distinguish the following cases.
[…]
@@
Follows from Lemma 12.27 together with the following observations. First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary 12.22. Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary 12.23. ∎
Proof of 2.

### v2

@@ 13. Vertex-transitive graphs
We may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$. By ˜9.41(U,W)(U,W) is a totally-nested tetra-separation. By Lemma˜13.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
We claim that $G[U]=K_{4}$. Assume otherwise, so $G[W]=K_{4}$. Then $G[W]=G_{s}$ for some bag $G_{s}\neq G_{t}$. Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma˜9.14. But then $\alpha(\mathcal{O})<4$ contradicts Lemma˜9.51.
Hence $G[U]=K_{4}$. Next, we claim that $S(U,W)$ contains no inward edge, and assume for a contradiction that $x_{1}\in S(U,W)$ is an inward edge with endvertex $u$ in the interior of $G_{t}$. By Lemma˜9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$. In particular, $y_{1}$ cannot be an outward edge. Since $S(U,W)$ consists of four edges and $y_{1}$ is one of them, $y_{1}$ must be an inward edge. Let $u$ be the endvertex of $y_{1}$ in the interior of $G_{t}$. Then $u\neq v$ by the matching-condition for $(U,W)$. Let $w$ be a neighbour of $x^{\prime}_{1}$ in $G$ but outside of $G_{t}$. Then the vertex-set $\{u,v,w\}$ is independent in $G$ and consists entirely of neighbours of $x^{\prime}_{1}$. Since $G[U]=K_{4}$ and $S(U,W)$ consists of four edges, $G$ is 4-regular and every vertex lies in a $K_{4}$ by vertex-transitivity. But $\{u,v,w\}$ witnesses that $x^{\prime}_{1}$ does not lie in a $K_{4}$, a contradiction. Therefore, $S(U,W)$ contains no inward edge.
@@
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$. Write $O=t(0)\,t(1)\ldots t(m)\,t(0)$ with indices in $\mathbb{Z}_{m+1}$. By Lemma˜13.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles. Then, by Lemma˜9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected. Hence, both of these torsos are good as they neighbour torsos that are 4-cycles. By induction, we get that $O$ has even length, that bags $G_{t(i)}$ are $K_{4}$’s for even $i$, and that torsos $H_{t(i)}$ are 4-cycles for odd $i$. By Lemma˜9.51, there are $\geqslant 4$ bags that are $K_{4}$’s. It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$. ∎
@@
Write $\mathcal{O}=:(O,\mathcal{G})$. By Lemma˜13.12, all bags of $\mathcal{O}$ are bad. Let $H_{t}$ be a torso of $\mathcal{O}$ on five vertices. Since the unique interior vertex $x$ of $H_{t}$ has degree 4, the graph $G$ is 4-regular. Since $H_{t}$ is bad, it is neighboured by a 3-connected torso $H_{s}$. Since $H_{s}$ is bad as well, it has $\leqslant 5$ vertices. That is, the torso of $H_{s}$ either is a $K_{4}$ or has exactly five vertices.
@@
It follows by induction that all torsos of $\mathcal{O}$ have exactly five vertices. By Lemma˜13.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag. By 4-regularity and since every bag looks like this, these are all the edges that there are in the bags.
Since each bag has a unique interior vertex, the set $X$ of all these vertices has size $|O|$. Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma˜9.51. If $|O|\geqslant 5$, then every vertex in $X$ is contained in exactly two induced 4-cycles in $G$, while every vertex in an adhesion set is contained in exactly one induced 4-cycle in $G$, contradicting vertex-transitivity. Therefore, we have $|O|=4$. Consider the graph H:=GXH:=G-X, which is either an 8-cycle or a disjoint union of two 4-cycles. If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in V(GX)V(G-X) is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity. Therefore $H$ is a disjoint union of two 4-cycles, and it follows that $G$ is the line graph of the cube. ∎
@@
If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma˜13.12, and it is of length $\geqslant 8$ by Lemma˜9.51. Hence, we may assume that all torsos of $\mathcal{O}$ are bad.
We claim that $\mathcal{O}$ has no 4-cycle as torso. Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma˜9.4, and they are good, a contradiction. Thus,
@@
We claim that if some torso of $\mathcal{O}$ is a triangle, then also its neighbouring torsos are triangles. For this, let $H_{t}$ and $H_{s}$ be neighbouring torsos with $H_{t}$ a triangle. By ˜8, the torso $H_{s}$ is either a triangle or a $K_{4}$. Assume for a contradiction that $H_{s}=K_{4}$. Let $\{u,v\}$ be the adhesion-set in between $H_{t}$ and $H_{s}$, so that $v$ is the tip of the triangle-torso $H_{t}$, say. Since every vertex has degree $\geqslant 4$, the edge $uv$ must exist in $G$, and $G$ must be 4-regular. Let $w$ be the vertex of the triangle $H_{t}$ besides $u$ and $v$. Let $G_{r}$ be the bag neighbouring $G_{t}$ besides $G_{s}$. Then $v$ has exactly one neighbour in the bag $G_{r}$. By ˜8, the torso $H_{r}$ can only be a triangle or a $K_{4}$. As $v$ has exactly one neighbour in the bag $G_{r}$, the torso $H_{r}$ cannot be a $K_{4}$, so it must be a triangle. But then the edge $vw$ exists by Lemma˜9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity. Therefore, every triangle-torso of $\mathcal{O}$ is neighboured by two triangle-torsos.
So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma˜9.4. Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma˜9.51 and since $G\neq K_{5}$.
So we may assume that $\mathcal{O}$ has no triangle-torsos. Then by ˜8, all torsos of $\mathcal{O}$ are $K_{4}$’s. Then either all adhesion-sets are spanned by edges in $G$, in which case $G$ is a cycle of $K_{4}$-bags; or none of them are, in which case $G$ is a cycle of $K_{2,2}$-bags. In both cases, the length is $\geqslant 4$ by Lemma˜9.51. ∎
@@
By the (Crossing Lemma). (4.1), all links have the same size 0, 11 or $2$. If all links have size 0, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma˜13.7. If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma˜13.10. Otherwise all links have size $2$, and we apply Lemma˜13.16. ∎
Observation 13.18.
Let $G^{\prime}$ be the clique-expansion of a graph $G$. Let $v$ be a vertex of degree $d$ in $G$, and let $C_{v}$ be the clique of $G^{\prime}$ that corresponds to $v$. Then every vertex in $C_{v}$ has degree $d$ in $G^{\prime}$.∎
Observation 13.19.
Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the clique-expansion of $G$. Then there is a unique partition $\mathcal{P}$ of $V(G^{\prime})$ into (the vertex-sets of) $|\mathcal{P}|=|V(G)|$ cliques. Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to GG.∎
Observation 13.20.
The clique-expansion of a graph $G$ is isomorphic to the graph $H$ with
$V(H)$ ${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
$E(H)$ ${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
@@
˜(i) $\Rightarrow$ ˜(ii). Let $(u_{1},v_{1}),(u_{2},v_{2})$ be ordered pairs of vertices in $G$ with $u_{i}v_{i}\in E(G)$. We will find an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. For this, we abbreviate $x_{i}:=(u_{i},v_{i})$. Let $P_{i}\in\mathcal{P}$ be the unique partition-class of $\mathcal{P}$ that contains $x_{i}$. Since $G^{\prime}$ is vertex-transitive by assumption, there is an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. Note that $\varphi$ then maps $P_{1}$ to $P_{2}$. We define $\varphi:=\gamma\circ\varphi^{\prime}\circ\gamma^{-1}$ by abuse of notation, using that $\operatorname{Aut}(G^{\prime})$ acts on $\mathcal{P}$. Clearly, $\varphi$ is a bijection $V(G)\to V(G)$. Moreover, we have
$\text{ }xy\in E(G)$
@@
˜(ii) $\Rightarrow$ ˜(i). We assume without loss of generality that $G^{\prime}$ takes the form described in ˜13.20. Let $x_{1},x_{2}\in V(G^{\prime})$ be two vertices. We will find an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. We write $x_{i}=:(u_{i},v_{i})$ using ˜13.20, so $u_{i}v_{i}$ is an edge of $G$. Since $G$ is arc-transitive, there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. Now $\varphi$ defines an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that sends every vertex $x=(u,v)$ to the vertex $\varphi^{\prime}(x):=(\varphi(u),\varphi(v))$. The automorphism $\varphi^{\prime}$ satisfies
@@
By Lemma˜13.21, it suffices to show that $G$ is not arc-transitive. Indeed, if $u$ is a vertex in $A$, then $G[N(u)]=K_{1}\sqcup K_{r-1}$ with $r-1\geqslant 2$. Let $v_{1}$ be the isolated vertex in $G[N(u)]$ and let $v_{2}\neq v_{1}$ be another vertex in $N(u)$. Then there is no automorphism of $G$ that maps $(u,v_{1})$ to $(u,v_{2})$. ∎
@@
Follows from Corollary˜13.22. ∎
@@
Let $G$ be a cycle of triangle-bags of length $\geqslant 6$, and let $\mathcal{O}$ be the corresponding cycle-decomposition. Note that every free edge of $\mathcal{O}$ is contained in exactly one triangle of $G$, while every other edge of $G$ is contained in exactly two triangles. So, $G$ is not edge-transitive. In particular, it is not arc-transitive. By Lemma˜13.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive. ∎
@@
This follows from Corollary˜13.17 together with the following observations. First, a cycle of $K_{4}$-bags of length $\geqslant 4$ is not 4-regular. Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary˜13.22. Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary˜13.24. ∎
@@
Let $G$ be a vertex-transitive essentially 5-connected finite graph. Then either $G$ is quasi-5-connected or there is a partition $\mathcal{P}$ of $V(G)$ into vertex-sets of $K_{4}$’s such that the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
@@
If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma˜2.3 (i). Hence, we may assume that $G$ has a tetra-separation $(A,B)$. Since $G$ is essentially 5-connected, the separator of $(A,B)$ consists of four edges and there is a side, say $A$, that induces a $K_{4}$ in $G$. Let $u\in A$ and observe that $G[N(u)]=K_{1}\sqcup K_{3}$; in particular, $u$ is contained in exactly one $K_{4}$. By vertex-transitivity, so is every vertex of $G$. It follows that $G$ admits a partition of its vertex-set into the vertex-sets of $K_{4}$’s. Moreover, $G$ is 4-regular since $u$ has degree 4 in $G$. Then it follows that the edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. ∎
@@
Let $G$ be a vertex-transitive essentially 5-connected finite graph. Then $G$ is one of the following:
@@
Assume that $G$ is not quasi-5-connected. Then by Lemma˜13.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. We distinguish the following cases.
Case 1: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are at least three $P_{1}$$P_{2}$ edges $e_{1},e_{2},e_{3}$ in $G$. Let $f_{i}\notin\{e_{1},e_{2},e_{3}\}$ be the remaining edge incident with a vertex in $P_{i}$. By 4-connectivity of $G$, $\{f_{1},f_{2}\}$ is not a mixed-2-separator. It follows that $f_{1}=f_{2}$, so $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length 4.
Case 2: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are exactly two $P_{1}$$P_{2}$ edges in $G$. Let $e_{1},e_{2}$ be the two edges incident with $P_{1}$ but not with $P_{2}$, and let $f_{1},f_{2}$ be the two edges incident with $P_{2}$ but not with $P_{1}$. We define $(A,B):=(P_{1}\cup P_{2},V(G)\smallsetminus(P_{1}\cup P_{2}))$. Then $S(A,B)=\{e_{1},e_{2},f_{1},f_{2}\}$, so $(A,B)$ is a tetra-separation. By assumption, there is a side of $(A,B)$ that induces a $K_{4}$. Since $|A|=8$, we must have $G[B]=K_{4}$. It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $6$.
Case 3: for every two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$, there is at most one $P_{1}$$P_{2}$ edge in $G$. Then we define $H:=G/\mathcal{P}$, which is a 4-regular graph. Moreover, observe that $G$ is the $K_{4}$-expansion of $H$. Then by Lemma˜13.21HH is arc-transitive. By Corollary˜13.2HH is 4-connected. ∎
@@
Follows from Lemma˜13.27 together with the following observations. First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary˜13.22. Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary˜13.23. ∎
Proposition 13.29.
A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
  1. (i)
    a quasi-5-connected vertex-transitive graph,
  2. (ii)
    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
  3. (iii)
    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
  4. (iv)
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
  5. (v)
    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
  1. (vi)
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure˜1.14),
  2. (vii)
    the line-graph of the cube (Figure˜1.15).
Proof.